An 70 kilogram guy descends from a height of 10.0 m by holding onto a rope that goes through a frictionless pulley to a 55 kg sandbag. What is the speed at which the man hits the ground if he starts from rest?


1
Expert's answer
2021-12-15T16:21:48-0500

"T-m_1g=-m_1a"


"T-m_2g=m_2a\to"


"a=\frac{m_1-m_2}{m_1+m_2}g=\frac{70-55}{70+55}\cdot9.8=1.18\ (m/s)"


"h=\frac{v^2}{2a}\to v=\sqrt{2ah}=\sqrt{2\cdot1.18\cdot10}=4.86\ (m/s)" . Answer